# Accurate Spatial Gene Expression Prediction by integrating Multi-resolution features

### v1

@@ Abstract
Recent advancements in Spatial Transcriptomics (ST) technology have facilitated detailed gene expression analysis within tissue contexts. However, the high costs and methodological limitations of ST necessitate a more robust predictive model. In response, this paper introduces TRIPLEX, a novel deep learning framework designed to predict spatial gene expression from Whole Slide Images (WSIs). TRIPLEX uniquely harnesses multi-resolution features, capturing cellular morphology at individual spots, the local context around these spots, and the global tissue organization. By integrating these features through an effective fusion strategy, TRIPLEX achieves accurate gene expression prediction. Our comprehensive benchmark study, conducted on three public ST datasets and supplemented with Visium data from 10X Genomics, demonstrates that TRIPLEX outperforms current state-of-the-art models in Mean Squared Error (MSE), Mean Absolute Error (MAE), and Pearson Correlation Coefficient (PCC). The model’s predictions align closely with ground truth gene expression profiles and tumor annotations, underscoring TRIPLEX’s potential in advancing cancer diagnosis and treatment. Code is available at https://github.com/NEXGEM/TRIPLEX
@@ 1 Introduction
Figure 1: Summary of our key idea to predict spatial gene expression. Left: We consider the features of a WSI in three different resolutions. Right: The gene expression profile to be predicted at the given spot.
@@
A number of studies have endeavored to address this challenge [8, 20, 29, 28, 26]. Approaches vary, with some predicting gene expression strictly from the tissue image confined within the spot’s boundaries [8], while others also take into account spatial dependencies between spot images [20, 29], or consider similarities to reference spots [28, 26]. However, we have noted several limitations inherent to these existing methodologies. Firstly, current methods primarily focus on spot images, neglecting the wealth of biological information available in the wider image context. By integrating both the specific spot and its surrounding environment, along with the holistic view of the entire histology image, we can access richer information, encompassing varied biological contexts. Second, models that consider interactions between spots [20, 29] face a limitation in processing the embedding of all patches in a WSI simultaneously. This approach, common in handling hundreds to thousands of patches within a WSI, limits the scalability of the patch embedding model due to resource constraints. Such limitations significantly impede the extraction of fine-grained, rich representations from each spot, thereby affecting the model’s ability to perform detailed analysis of WSIs. Thirdly, model performance is frequently overestimated because of inadequate validation, such as using the limited size of dataset [8] sometimes without cross-validation [26] and training/testing with replicates from the same patient [20, 29, 28]. The limited size of ST datasets means that exclusive reliance on a single dataset for model evaluation can hinder an accurate assessment of the model’s capabilities, thereby emphasizing the necessity for cross-validation. The issue is compounded when replicate data from the same patient, often featuring nearly identical image-gene expression pairs, are used in both training and testing phases. This can lead to an inflated perception of a model’s effectiveness, as it may not accurately reflect the model’s ability to generalize to new, unseen data. Lastly, the use of disparate datasets, diverse normalization methods, and varied evaluation techniques in existing research studies compounds the challenge of conducting fair comparisons of the models.
@@
To overcome these limitations, we present TRIPLEX, an innovative deep learning framework designed to leverage multi-resolution features from histology images for robustly predicting spatial gene expression levels. TRIPLEX extracts three distinct types of features corresponding to different resolutions: the target spot image (representing the specific spot, whose gene expression to be predicted), the neighbor view (encompassing a wider area around the spot), and the global view (comprising the aggregate of all spot images). These features capture varying levels of biological information—ranging from the detailed cell morphology in the target spot image, to the surrounding tissue phenotype, and the overall tissue microenvironment in the WSI, as illustrated in Figure 1. Each is integral to understanding the spatial gene expression levels of the given spot. TRIPLEX employs separate encoders to extract these features from WSIs, each focusing on its assigned resolution to efficiently capture relevant details. These features are then integrated via a fusion layer for effective gene expression prediction. This approach allows TRIPLEX to utilize resolution-specific information, thereby enhancing prediction accuracy while avoiding significant increases in computational costs.
@@
Our study sets a new benchmark in spatial gene expression prediction, comparing our model, TRIPLEX, against five prior studies [8, 20, 29, 28, 26] under uniform experimental conditions. We conduct internal evaluations using three public Spatial Transcriptomics (ST) datasets [1, 11, 8] and external validations using higher-resolution Visium data from 10X Genomics. Our validation procedure strictly avoids mixing patient sample replicates between training and testing datasets, a significant departure from previous methods [20, 29, 28], and employs rigorous cross-validation. Our results indicate that TRIPLEX surpasses existing models in terms of Mean Squared Error (MSE), Mean Absolute Error (MAE), and Pearson Correlation Coefficient (PCC) in both internal and external evaluations. Furthermore, we provide visualizations of the expression distributions for specific genes commonly associated with cancer. These visualizations reveal that our model’s predictions align more closely with actual gene expression data and tumor annotations, demonstrating its enhanced predictive accuracy.
@@ 2 Related Work
In this section, we delve into studies pertinent to our research. For clarity, throughout this discussion, the term ’spot’ will be used to denote a predefined unit region within a WSI where gene expression is quantified. Moreover, we will use ’target spot’ to specifically refer to the spot within a WSI for which we seek to predict gene expression.
Spatial gene expression prediction from WSIs via deep learning We review the pioneering works in this field, which aim to predict spatial gene expression from WSIs. ST-Net [8] utilizes a standard transfer learning strategy, training a Densenet121 model [9]—pretrained on ImageNet—using histology images as input and gene expression as labels. Following this, HisToGene [20] leverages Vision Transformers (ViT) [7] to account for correlations among patches in a WSI, thereby predicting gene expression from global-context aware features. Further developing this concept, Hist2ST [29] enhances the approach by emphasizing patch embedding using ConvMixer [25] and aggregating neighborhood information through graph convolution network [15]. While these previous works share similarities with our methodology, they predominantly process patch and global embedding sequentially, often overlooking the neighboring information around the target spot. In contrast, our method sets itself apart by concurrently extracting critical features at three distinct resolutions, including the neighbor view, and integrating them for gene expression prediction. In a different vein, EGN [28] adopts exemplar learning for predicting gene expression from histology images. This method dynamically selects the most analogous exemplars from a target spot within a WSI to enhance prediction accuracy. Additionally, BLEEP [26] introduces a bi-modal embedding framework similar to CLIP [21] to co-embed spot images and gene expression. After training, this model imputes the gene expression of a query spot using the gene expression set from a reference dataset.
@@ 4.4 Ablation study
In this section, we demonstrate the contributions of each method to gene expression prediction through an ablation study of our proposed model. Specifically, we aim to observe the contributions of the three modules in our model and investigate the extent to which the position encoding generator and our proposed fusion approach contribute to gene expression prediction. Here we only present experimental results on the SCC dataset, but we observe similar outcomes on other datasets as well. Additional experiment results can be found in supplementary material.
@@
Position Encoding Generator (PEG) We further compare our APEG to different methods of positional encoding, including (1) without PEG and (2) with PEG [5], as shown in Table 4. In the conventional PEG implementation, we added zero padding to the global token to make the number of tokens a perfect square, followed by the standard squaring procedure. The results indicate that APEG not only outperforms the model without any PEG but also surpasses the conventional PEG. This implies that APEG more effectively encodes spatial distribution within the global tokens, enhancing the overall model performance.
MSE MAE PCC(M) PCC(T)
w/o PEG 0.280 0.413 0.360 0.480
PEG [5] 0.276 0.411 0.364 0.479
APEG 0.268 0.404 0.374 0.490
@@
Fusion Method and Fusion Loss We also investigated the impact of different fusion methods for multi-resolution features, comparing our model’s fusion layer with traditional feature fusion techniques like summation, concatenation, and attentional pooling [10]. Additionally, we assessed the contribution of fusion loss to the model’s performance. As shown in Table 5, integrating multi-resolution features using our fusion layer yields superior results compared to these conventional techniques. Moreover, the marked improvement in performance when incorporating fusion loss indicates its effectiveness in integrating the three types of features. This result highlights the significance of our fusion approach in achieving high accuracy in gene expression prediction.
Fusion method MSE MAE PCC(M) PCC(T)
Summation 0.297 0.425 0.341 0.464
Concatenation 0.293 0.422 0.348 0.474
Attentional pooling 0.293 0.423 0.353 0.473
fusion layer 0.268 0.404 0.374 0.490
Fusion loss MSE MAE PCC(M) PCC(T)
w/o fusion loss 0.292 0.423 0.358 0.469
w/ fusion loss 0.268 0.404 0.374 0.490
@@ 5 Conclusion
A Evaluation methodology for multi-output regression problem
@@
B Method Details
B.1 Processing of input data
@@
B.2 Method details for Target Encoder
@@
B.3 Method details for Neighbor Encoder
The neighbor encoder is designed to embed local information surrounding the target spot. It processes features extracted from 25 images, each of size 224x224, representing the neighbor view. This approach closely aligns with the Vision Transformer (ViT) [7] architecture, incorporating self-attention mechanisms with relative position encoding and fully-connected layers applied to the input tokens. A key deviation from the standard ViT model is the exclusion of the patch embedding module for 2D images. Instead, we directly utilize pre-extracted features as our input values. Details regarding specific hyperparameters and their settings will be discussed in the forthcoming section, ”Additional Implementation Details and Experimental Results.” In a manner similar to the target layer, an additional fully-connected layer is attached to the pooled token of the neighbor tokens. This layer is also specifically tasked with the prediction of gene expression.
B.4 Method details for Global Encoder
@@
B.5 Method details for Fusion Layer
@@
B.6 Implementation Details for Baselines
@@
BLEEP BLEEP [26] is a bi-modal learning model designed to co-embed histology images and gene expression levels. In our implementation, we adhere to the core architecture of BLEEP, with minor hyperparameter adjustments as detailed in Section 3.6. For inference, BLEEP utilizes the top-k nearest gene expression levels to predict gene expression for query spot images. Among the three methods suggested - ”simple” (using the top 1 value), ”simple average” (averaging the top k values), and ”weighted average” (using a weighted average of the top k values) - we employ the ”simple average” approach, specifically using the top-50 nearest gene expression levels. This choice was based on its superior performance in our tests. Additionally, we opted not to use Harmony [16] for batch correction of gene expression levels, as Harmony is geared towards correcting PCA embeddings rather than raw gene expression levels. Given our dataset’s limited slide range (12 to 68), such batch correction might inadvertently reduce the training data’s diversity, thereby increasing the risk of overfitting.
@@
C Additional implementation details and experimental results
C.1 Description of datasets
We evaluate our model on three distinct datasets: BC1 and BC2 (breast cancer datasets) and SCC (a skin cancer dataset). We focus on the top 250 genes with high gene expression levels in each dataset as labels for prediction. Furthermore, we calculate the average ranks of well-predicted genes during cross-validation to identify the top 50 ”highly predictive genes.” These genes are then used to compute the PCC (H). For detailed summaries of each dataset and the specific genes selected, please refer to Figures 6, 7, and 8.
C.2 Implementation Details for Experiments
@@
C.3 Implementation Details for Ablation Studies
@@
C.4 Computational Cost Comparison
@@
C.5 Comparison of MAE in the cross-validation experiments
@@
C.6 Contribution of the Neighbor View size
We examine how the performance of TRIPLEX is influenced by the expansion of the neighbor view size. Here, the term ”number of neighbors” refers to the count of 224x224 patches along an axis. Results are illustrated in Figure 10. When evaluated in terms of MES and PCC, it is observed that enlarging the neighbor view size does not consistently result in performance gains. In fact, as the number of neighbors increases, a decrease in performance is noted. This pattern suggests that 1,120x1,120 sized neighbor view (number of neighbors: 5) is an efficient configuration, striking a balance between capturing detailed neighboring information relevant to the target and maintaining manageable computational costs.
C.7 Performance Discrepancy Between Our Experimental Results and Existing Implementations
@@
C.8 Additional Ablation Studies
@@
C.9 Detailed Hyperparameter Settings
@@
Figure 10: The performance varies with the size of the neighbor view. Variations in MSE (Left) and PCC (M) (Right) relative to the size. ”Number of neighbors” represents the count of 224x224 patches along an axis
D Additional Visualizations
In this section, we present additional visualizations focusing on the spatial expression distribution prediction of the GNAS gene, as shown in Figures 11 to 14. These visualizations include four additional samples from the BC1 dataset and 20 samples from the BC2 dataset. In analyzing these visualizations, we observe a high degree of consistency between the GNAS expression distribution and the annotations provided by pathologists. Notably, the predictions made by TRIPLEX demonstrate a markedly high accuracy, as quantitatively assessed against benchmark metrics, and align closely with tumor annotations. This consistency is particularly evident when compared to the predictions from other baseline models [8, 20, 29, 28, 26].

### v2

@@ Abstract
Recent advancements in Spatial Transcriptomics (ST) technology have facilitated detailed gene expression analysis within tissue contexts. However, the high costs and methodological limitations of ST necessitate a more robust predictive model. In response, this paper introduces TRIPLEX, a novel deep learning framework designed to predict spatial gene expression from Whole Slide Images (WSIs). TRIPLEX uniquely harnesses multi-resolution features, capturing cellular morphology at individual spots, the local context around these spots, and the global tissue organization. By integrating these features through an effective fusion strategy, TRIPLEX achieves accurate gene expression prediction. Our comprehensive benchmark study, conducted on three public ST datasets and supplemented with Visium data from 10X Genomics, demonstrates that TRIPLEX outperforms current state-of-the-art models in Mean Squared Error (MSE), Mean Absolute Error (MAE), and Pearson Correlation Coefficient (PCC). The model’s predictions align closely with ground truth gene expression profiles and tumor annotations, underscoring TRIPLEX’s potential in advancing cancer diagnosis and treatment.
@@ 1 Introduction
A number of studies have endeavored to address this challenge [8, 20, 29, 28, 26]. Approaches vary, with some predicting gene expression strictly from the tissue image confined within the spot’s boundaries [8], while others also take into account spatial dependencies between spot images [20, 29], or consider similarities to reference spots [28, 26]. However, we have noted several limitations inherent to these existing methodologies. Firstly, current methods primarily focus on spot images, neglecting the wealth of biological information available in the wider image context. By integrating both the specific spot and its surrounding environment, along with the holistic view of the entire histology image, we can access richer information, encompassing varied biological contexts. Secondly, models that consider interactions between spots [20, 29] face a limitation in processing the embedding of all patches in a WSI simultaneously. This approach, common in handling hundreds to thousands of patches within a WSI, limits the scalability of the patch embedding model due to resource constraints. Such limitations significantly impede the extraction of fine-grained, rich representations from each spot, thereby affecting the model’s ability to perform detailed analysis of WSIs. Thirdly, model performance is frequently overestimated because of inadequate validation, such as using the limited size of dataset [8] sometimes without cross-validation [26] and training/testing with replicates from the same patient [20, 29, 28]. The limited size of ST datasets means that exclusive reliance on a single dataset for model evaluation can hinder an accurate assessment of the model’s capabilities, thereby emphasizing the necessity for cross-validation. The issue is compounded when replicate data from the same patient, often featuring nearly identical image-gene expression pairs, are used in both training and testing phases. This can lead to an inflated perception of a model’s effectiveness, as it may not accurately reflect the model’s ability to generalize to new, unseen data. Lastly, the use of disparate datasets, diverse normalization methods, and varied evaluation techniques in existing research studies compounds the challenge of conducting fair comparisons of the models.
@@
Addressing these limitations, we present TRIPLEX, an innovative deep learning framework designed to leverage multi-resolution features from histology images for robustly predicting spatial gene expression levels. TRIPLEX extracts three distinct types of features corresponding to different resolutions: the target spot image (representing the specific spot, whose gene expression to be predicted), the neighbor view (encompassing a wider area around the spot), and the global view (comprising the aggregate of all spot images). These features capture varying levels of biological information—ranging from the detailed cell morphology in the target spot image, to the surrounding tissue phenotype, and the overall tissue microenvironment in the WSI. Each is integral to understanding the spatial gene expression levels of the given spot. TRIPLEX employs separate encoders to extract these features from WSIs, each focusing on its assigned resolution to efficiently capture relevant details. For neighbour or global view with larger resolution, pre-extracted features are used to reduce the burden of computational cost, while for target spot images, encoders are fully updated to extract fine-grained information. These features are then integrated via a fusion layer for effective gene expression prediction. This approach allows TRIPLEX to utilize resolution-specific information, thereby enhancing prediction accuracy while avoiding significant increases in computational costs.
@@
Our study sets a new benchmark in spatial gene expression prediction, comparing our model, TRIPLEX, against five prior studies [8, 20, 29, 28, 26] under uniform experimental conditions. We conduct internal evaluations using three public Spatial Transcriptomics (ST) datasets [1, 11, 8] and external validations using higher-resolution Visium data from 10X Genomics. Our validation procedure strictly avoids mixing patient sample replicates between training and testing datasets, a significant departure from previous methods [20, 29, 28], and employs rigorous cross-validation. Our results indicate that TRIPLEX surpasses existing models in terms of Mean Squared Error (MSE), Mean Absolute Error (MAE), and Pearson Correlation Coefficient (PCC) in both internal and external evaluations. Furthermore, we provide visualizations of the expression distributions for a specific gene commonly associated with cancer. These visualizations reveal that our model’s predictions align more closely with actual gene expression data and tumor annotations, demonstrating its enhanced predictive accuracy.
@@ 2 Related Work
In this section, we delve into studies pertinent to our research. For clarity, the term ’spot’ will be used to denote a predefined unit region within a WSI where gene expression is quantified. Moreover, we will use ’target spot’ to specifically refer to the spot within a WSI for which we seek to predict gene expression.
Spatial gene expression prediction from WSIs via deep learning We review the pioneering works in this field, which aim to predict spatial gene expression from WSIs. ST-Net [8] utilizes a standard transfer learning strategy, training a Densenet121 model [9]—pretrained on ImageNet—using histology images as input and gene expression as labels. Following this, HisToGene [20] leverages Vision Transformers (ViT) [7] to account for correlations among patches in a WSI, thereby predicting gene expression from global-context aware features. Further developing this concept, Hist2ST [29] enhances the approach by emphasizing patch embedding using ConvMixer [25] and aggregating neighborhood information through graph convolution network [15]. While these previous works [20, 29] share similarities with our methodology, they predominantly process patch and global embedding sequentially, often overlooking the neighboring information around the target spot. In contrast, our method sets itself apart by concurrently extracting critical features at three distinct resolutions, including the neighbor view, and integrating them for gene expression prediction. In a different vein, EGN [28] adopts exemplar learning for predicting gene expression from histology images. This method dynamically selects the most analogous exemplars from a target spot within a WSI to enhance prediction accuracy. Additionally, BLEEP [26] introduces a bi-modal embedding framework similar to CLIP [21] to co-embed spot images and gene expression. After training, this model imputes the gene expression of a query spot using the retrieved gene expression set from a reference dataset.
@@ 4.4 Ablation study
In this section, we demonstrate the contributions of each method to gene expression prediction through an ablation study of our proposed model. Specifically, we aim to observe the contributions of the three modules in our model and investigate the extent to which the position encoding generator and our proposed fusion approach contribute to gene expression prediction. Here we only present experimental results on the SCC dataset, but we have observed similar outcomes on other datasets as well. Additional experiment results can be found in supplementary material.
@@
Position Encoding Generator (PEG) We further compare our APEG to different methods of positional encoding, including (1) without PEG and (2) with PEG [5], as shown in Table 4. In the conventional PEG implementation, we add zero padding to the global token to make the number of tokens a perfect square, followed by the standard squaring procedure. The results indicate that APEG not only outperforms the model without any PEG but also surpasses the conventional PEG. This implies that APEG more effectively encodes spatial distribution within the global tokens, enhancing the overall model performance.
MSE MAE PCC(M) PCC(T) w/o PEG 0.280 0.413 0.360 0.480 PEG [5] 0.276 0.411 0.364 0.479 APEG 0.268 0.404 0.374 0.490
@@
Fusion Method and Fusion Loss We also investigate the impact of different fusion methods for multi-resolution features, comparing our model’s fusion layer with traditional feature fusion techniques like summation, concatenation, and attentional pooling [10]. Additionally, we assess the contribution of fusion loss to the model’s performance. As shown in Table 5, integrating multi-resolution features using our fusion layer yields superior results compared to these conventional techniques. Moreover, the marked improvement in performance when incorporating fusion loss indicates its effectiveness in integrating the three types of features. This result highlights the significance of our fusion approach in achieving high accuracy in gene expression prediction.
Fusion method MSE MAE PCC(M) PCC(T) Summation 0.297 0.425 0.341 0.464 Concatenation 0.293 0.422 0.348 0.474 Attentional pooling 0.293 0.423 0.353 0.473 fusion layer 0.268 0.404 0.374 0.490 Fusion loss MSE MAE PCC(M) PCC(T) w/o fusion loss 0.292 0.423 0.358 0.469 w/ fusion loss 0.268 0.404 0.374 0.490
@@ 5 Conclusion
Supplementary Material
F Evaluation methodology for multi-output regression problem
@@
G Method Details
G.1 Processing of input data
@@
G.2 Method details for Target Encoder
@@
G.3 Method details for Neighbor Encoder
The neighbor encoder is designed to embed local information surrounding the target spot. It processes features extracted from 25 images, each of size 224x224, representing the neighbor view. This approach closely aligns with the Vision Transformer (ViT) [7] architecture, incorporating self-attention mechanisms with relative position encoding and fully-connected layers applied to the input tokens. A key deviation from the standard ViT model is the exclusion of the patch embedding module for 2D images. Instead, we directly utilize pre-extracted features as our input values. Details regarding specific hyperparameters and their settings will be discussed in the forthcoming section, "Additional Implementation Details and Experimental Results." In a manner similar to the target layer, an additional fully-connected layer is attached to the pooled token of the neighbor tokens. This layer is also specifically tasked with the prediction of gene expression.
G.4 Method details for Global Encoder
@@
G.5 Method details for Fusion Layer
@@
G.6 Implementation Details for Baselines
@@
BLEEP BLEEP [26] is a bi-modal learning model designed to co-embed histology images and gene expression levels. In our implementation, we adhere to the core architecture of BLEEP, with minor hyperparameter adjustments as detailed in Section 3.6. For inference, BLEEP utilizes the top-k nearest gene expression levels to predict gene expression for query spot images. Among the three methods suggested - "simple" (using the top 1 value), "simple average" (averaging the top k values), and "weighted average" (using a weighted average of the top k values) - we employ the "simple average" approach, specifically using the top-50 nearest gene expression levels. This choice was based on its superior performance in our tests. Additionally, we opted not to use Harmony [16] for batch correction of gene expression levels, as Harmony is geared towards correcting PCA embeddings rather than raw gene expression levels. Given our dataset’s limited slide range (12 to 68), such batch correction might inadvertently reduce the training data’s diversity, thereby increasing the risk of overfitting.
@@
H Additional implementation details and experimental results
H.1 Description of datasets
We evaluate our model on three distinct datasets: BC1 and BC2 (breast cancer datasets) and SCC (a skin cancer dataset). We focus on the top 250 genes with high gene expression levels in each dataset as labels for prediction. Furthermore, we calculate the average ranks of well-predicted genes during cross-validation to identify the top 50 "highly predictive genes." These genes are then used to compute the PCC (H). For detailed summaries of each dataset and the specific genes selected, please refer to Figures 5, 6, and 7.
H.2 Implementation Details for Experiments
@@
H.3 Implementation Details for Ablation Studies
@@
H.4 Computational Cost Comparison
@@
H.5 Comparison of MAE in the cross-validation experiments
@@
H.6 Contribution of the Neighbor View size
We examine how the performance of TRIPLEX is influenced by the expansion of the neighbor view size. Here, the term "number of neighbors" refers to the count of 224x224 patches along an axis. Results are illustrated in Figure 9. When evaluated in terms of MES and PCC, it is observed that enlarging the neighbor view size does not consistently result in performance gains. In fact, as the number of neighbors increases, a decrease in performance is noted. This pattern suggests that 1,120x1,120 sized neighbor view (number of neighbors: 5) is an efficient configuration, striking a balance between capturing detailed neighboring information relevant to the target and maintaining manageable computational costs.
H.7 Performance Discrepancy Between Our Experimental Results and Existing Implementations
@@
H.8 Additional Ablation Studies
@@
H.9 Detailed Hyperparameter Settings
@@
Figure 9: The performance varies with the size of the neighbor view. Variations in MSE (Left) and PCC (M) (Right) relative to the size. "Number of neighbors" represents the count of 224x224 patches along an axis
I Additional Visualizations
In this section, we present additional visualizations focusing on the spatial expression distribution prediction of the GNAS gene, as shown in Figures 10 to 13. These visualizations include four additional samples from the BC1 dataset and 20 samples from the BC2 dataset. In analyzing these visualizations, we observe a high degree of consistency between the GNAS expression distribution and the annotations provided by pathologists. Notably, the predictions made by TRIPLEX demonstrate a markedly high accuracy, as quantitatively assessed against benchmark metrics, and align closely with tumor annotations. This consistency is particularly evident when compared to the predictions from other baseline models [8, 20, 29, 28, 26].